Laser light (488 nm) passes through double slits forming a series of maxima on a screen 4.00 m away. If the distance between the central maximum and the first order maximum is 1.5 cm, what is the slit separation?

1.63x10^-4 m, This is the currect answer i just do not know how to reach it.
Answer #1

dsinθ = mλ

d(Δy/D) = mλ

d*(0.015/4)=1*488*10^-9

d= 1.39*10^-4 m

This is the correct answer. Please check your answer.